# Word length predicts word order: “Min-max”-ing drives language evolution

### v1

@@ 3.1 A “Min-Max” theory of processing
Another observation from the results reported here is that information content seems to trump length. We find that long nouns are used with more frequency in SV languages, while long verbs are used with more frequency in VS languages. This indicates that (maximizing) information content is more important to speakers than (minimizing) processing complexity. This could explain the “agent/subject-first preference” in experiments (Sauppe et al. 2023; Culbertson et al. 2012), as the “information maximization” pressure wins out in cases where this most efficiently meets communicative goals.
@@
The idea that speakers balance multiple considerations in production is not new (Pike 1967; Jackendoff & Pinker 2005; Seržant & Moroz 2022; Hörberg & Sjons 2023) but what is novel is the idea that all humans experience a specific universal pressure to balance production/processing efficiency with information content. This pressure is observable crosslinguistically, and would apply at multiple processing levels, regardless of whether a speaker is accessing a universal grammar (“innate”) or inferring structure on the basis of the input they receive (“usage”). Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers. This study shows that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically. The pressures identified here have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
3.2 Limitations
@@
One limitation in this study is the use of automated processes to identify part of speech tags in the data. Ideally, all the data should be hand-annotated for parts of speech, which would allow for better comparison of the different word categories across languages. Given the time consuming nature of this task and the need for expert annotators, this is not yet possible, but ongoing research could address this by sourcing experts to annotate the data. Although I have shown (in Ring 2025 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.
@@
Another limitation is that the dataset used, while exceedingly large in comparison to previous efforts, still only represents about a quarter of the world’s languages. Since the majority of the world’s languages are not represented here, it is not clear that these findings can be fully generalized to all languages of the world, though I have shown in section 2.4 that much can be learned from sampling. Additionally, the contrast between intransitive SV/VS/free word order is somewhat of an oversimplification of word order patterns. Given the current obstacles to annotation, it is a necessary simplification for this study, however future work could expand to a more diverse set of word/constituent orders and potential relationships with word length and/or complexity. Given the nature of the dataset used, it is also possible to pursue a gradient approach to word order, as advocated by Levshina et al. (2023). The issue of word complexity also could use further expansion to cover a broader range of possibilities than simply the length of a word based on a set of (roman) characters.
@@
A final note is that our conclusions are only as good as the data that we have available to us. The tagged PBC is a significant step in the right direction, toward a representative sample of large tagged corpora for languages of the world. Systematic investigation of this dataset can give us insight into language typology (the current state of language), as well as pressures that impact this realization in diachrony, both in terms of how languages have developed (their history) and how they might develop (their future). On the basis of this investigation I have proposed a parsimonious “Min-Max” theory of processing and information structure that underlies language evolution, whereby speakers minimize processing effort while maximizing information gain. This proposal is in keeping with at least one existing information-theoretic proposal (IB) as well as theories of processing (PGCH; surprisal) and can explain the correlation between the length of nouns/verbs and basic word order, given that these features allow us to predict the word order of a language. It is as yet unclear whether such a principle would apply at multiple hierarchical levels within language, which is an area that deserves further careful investigation with datasets such as the tagged PBC. For now, I have demonstrated that noun/verb lengths are more explanatory than descent or geography and that for languages with corpus data, word length predicts word order.

### v2

@@ 3.1 A “Min-Max” theory of processing
Another observation from the results reported here is that information content seems to trump length. We find that long nouns are used with more frequency in SV languages, while long verbs are used with more frequency in VS languages. This indicates that (maximizing) information content is more important to speakers than (minimizing) processing complexity. This could explain the “agent/subject-first preference” in experiments (Culbertson et al. 2012; Sauppe et al. 2023), as the “information maximization” pressure wins out in cases where this most efficiently meets communicative goals.
@@
The idea that speakers balance multiple considerations in production is not new (Pike 1967; Jackendoff & Pinker 2005; Seržant & Moroz 2022; Hörberg & Sjons 2023) but what is novel is the idea that all humans experience a specific universal pressure to balance production/processing efficiency with information content. This pressure is observable crosslinguistically, and would apply at multiple processing levels, regardless of whether a speaker is accessing a universal grammar (“innate”) or inferring structure on the basis of the input they receive (“usage”). Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers. The studies presented here show that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically. The pressures identified in this paper have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
3.2 Limitations and future directions
@@
One limitation in this set of studies is the use of automated processes to identify part of speech tags in the data. Ideally, all the data should be hand-annotated for parts of speech, which would allow for better comparison of the different word categories across languages. Given the time consuming nature of this task and the need for expert annotators, this is not yet possible, but ongoing research could address this by sourcing experts to annotate the data. Although I have shown (in Ring 2025 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.
@@
Another limitation is that the dataset used, while exceedingly large in comparison to previous efforts, still only represents about a quarter of the world’s languages. Since the majority of the world’s languages are not represented here, it is not clear that these findings can be fully generalized to all languages of the world, though I have shown in section 2.3 that much can be learned from sampling. Additionally, the contrast between intransitive SV/VS/free word order is somewhat of an oversimplification of word order patterns. Given the current obstacles to annotation, it is a necessary simplification for this study, however future work could expand to a more diverse set of word/constituent orders and potential relationships with word length and/or complexity. Given the nature of the dataset used, it is also possible to pursue a gradient approach to word order, as advocated by Levshina et al. (2023). The issue of word complexity could additionally use further expansion to cover a broader range of possibilities than simply the length of a word based on a set of (roman) characters.
@@
A final note is that our conclusions are only as good as the data that we have available to us. The taggedPBC is a significant step in the right direction, toward a representative sample of large tagged corpora for languages of the world. Systematic investigation of this dataset can give us insight into language typology (the current state of language), as well as pressures that impact this realization in diachrony, both in terms of how languages have developed (their history) and how they might develop (their future). On the basis of this investigation I have proposed a parsimonious “Min-Max” theory of processing and information structure that underlies language evolution, whereby speakers minimize processing effort while maximizing information gain. This proposal is in keeping with at least one existing information-theoretic proposal (IB) as well as theories of processing (efficiency; PGCH; surprisal) and can explain the correlation between the length of nouns/verbs and basic word order, given that these features allow us to predict the word order of a language. It is as yet unclear whether such a principle would apply at multiple hierarchical levels within language, which is an area that deserves further careful investigation with datasets such as the taggedPBC. For now, I have demonstrated that noun/verb lengths are more explanatory than descent or geography and that for languages with corpus data, word length predicts word order.